OpenStudy (anonymous):

Solve: 2n(2n + 2)=10+2n (squared)

OpenStudy (bibby):

distribute 2n to (2n+2) first

OpenStudy (anonymous):

I did it step by step but I didn't get the right answer surprisingly...

OpenStudy (bibby):

\(2n(2n + 2)=10+2n^2\) ?

OpenStudy (anonymous):

Yes

OpenStudy (xmissalycatx):

What you have to do first is distribute 2n to everything in the parentheses behind it, so you would get 4n(squared) + 4n = 10 + 2n(squared).

OpenStudy (xmissalycatx):

Next you would subtract 2n2 from both sides so that your value of 10 is by itself.. Now you have 2n^2 + 2n = 10

OpenStudy (anonymous):

What answer do you get?

OpenStudy (bibby):

\(2n(2n + 2)=10+2n^2\\4n^2 + 4n=10+2n^2\\2n^2+4n-10=0\\2(n^2+2n-5)=0\\n^2+2n-5=0\\n=-1 \pm \sqrt 6\)